Help me solve this for t please.
F = tu^2 + tvu for t.

Answer by stanbon(75887)

Help me solve this for t please.
F = tu^2 + tvu for t.
------
Factor out the "t":
t(u^2 + vu) = F
-------
t = F/(u^2+vu)
------------------
Cheers,
Stan H.